OpenStudy (anonymous):

What is the value for x in the proportion The quantity of x plus 8 all over the quantity of 5x minus 2 equals 3 over 8?

jimthompson5910 (jim_thompson5910):

the problem is \[\large \frac{x+8}{5x-2} = \frac{3}{8}\] right?

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

\[\large \frac{x+8}{5x-2} = \frac{3}{8}\] \[\large 8(x+8) = 3(5x-2)\] \[\large 8x+64 = 15x-6\] I'll let you finish up. Tell me what answer you get so I can check

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

x=10?

jimthompson5910 (jim_thompson5910):

you got it, nice work

OpenStudy (anonymous):

yay! thank u(:

jimthompson5910 (jim_thompson5910):

you're welcome